# This work has been submitted to the IEEE Access for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible. xx.xxxx/ACCESS.2024.xxxxxxx

### v2

@@ III-C2 Weight Quantization
Learning MNIST-DVS Lava
[332] [333]
DSD-SNN Task-IL, Supervised LIF - Back- - - Permuted MNIST BrainCog
[316],(2023) Class-IL Propagation Not-MNIST [330], [334]
Split-CIFAR100
SOR-SNN Class-IL Supervised LIF - Back- Adaptive - CIFAR100 [335], -
[317],(2023) Propagation ImageNet [336]
[318],(2023) Class-IL, Supervised LIF BPTT Fixed MNIST PyTorch
TA Fixed [337]
[292],(2023) Sample-IL, LIF - BPTT - Time SHD
Class-IL Supervised Compression [338]
Order R-STDP $a^{-}>-0.1125$ EMNIST [339] [340]
[320],(2024) DIL CIFAR100
III-D Summary of NCL
@@
The state-of-the-art methods have performed initial studies for enabling NCL (Table IV). We observe that methods with bio-plausible learning rules (e.g., STDP) facilitate unsupervised learning, which is suitable for OCL scenario. These works mainly employ fully-connected SNN architectures (see Table V), consisting of input layer and a pair of excitatory-inhibitory layer. As the input data grows in complexity, fully-connected SNNs struggle to capture the important hierarchical features without increasing the network size. Larger architectures often lead to better accuracy but require larger storage sizes, which may exceed the resource constraints. This limits the deployable networks on practical hardware platforms, and make their performances limited to simple datasets (e.g., MNIST) with simple CL scenarios (e.g., Split-MNIST). Meanwhile, methods that employ supervised-based learning rules (e.g., surrogate gradient-based backpropagation) can achieve acceptable performance on complex datasets (e.g., ImageNet). However, they are not suitable for OCL scenario, as they require costly labeled data and power-hungry training. Furthermore, these works mainly focus on energy-efficiency, and have not optimized resource utilization. Therefore, alternative methods are required especially for enabling energy-efficient SNNs with OCL capabilities, which are beneficial for many real-world application use-cases.
@@
TABLE V: Comparative analysis of the state-of-the-art in terms of network architecture, model size, performance, memory footprint, and power/energy savings. Note, FC: fully-connected layer, and CONV: convolutional layer.
@@
Work SNN Model Performance Memory Power/Energy
Architecture Size Footprint Consumption
ASP [97] 2 FC layers 200 excitatory & Achieved avg. accuracy of 75.8% (exp. decay) n/a n/a
(1 input layer & 200 inhibitory and 74.5% (linear decay), outperforming STDP
1 excit.-inhib. layer) neurons by $\sim$5.7%, which had 70.1% accuracy
SpikeDyn 2 FC layers (1 input 200 and 400 Improved accuracy by 21% for the last learned $<$3000 KB Avg. reduction of 51%
[41] & 1 excit. layers) excit. neurons task and 8% for previous tasks than ASP (training) and 37%
@@
2048, 1024 neurons for Split-MNIST and 74% for MNIST-DVS
per-layer
@@
[289] 3 CONV layers 30-250-200 Achieved 92.0±0.1% accuracy for digits in n/a n/a
@@ IV Real-World Application Use-Cases
In this section, we discuss the real-world application use-cases that benefit from bio-plausible SNNs with OCL capabilities. Additionally, we also highlight the current state-of-the-art for these application use-cases.
@@ IV-A1 Object Recognition for Humanoid Robots
Adaptive Robots need to identify, classify, and locate objects within an image/video. For this, Neural Engineering Framework (NEF) is explored for developing Neural State Machine (NSM) framework leveraging SNNs, managing the allocation of new neurons (Fig. 26), where data may be collected on-the-fly without any labels [341]. However, it relies on supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, thus it is not suitable for OCL scenario.
@@
IV-A2 Inverse Kinematics and PID Control
Recent study explored Neural Engineering Framework (NEF) for developing neuromorphic algorithms for Inverse Kinematics (IK) and Proportional-Integral-Derivative (PID) control systems, supporting online learning of control signals, using Prescribed Error Sensitivity (PES) [342]. Fig. 27(a) shows the model schematic for neuromorphic IK with online learning, whose target is to efficiently compute the joint angles required to place the robots’ end-effector at the desired position. Meanwhile, the PID controller aims to ensure that the robots’ actuators accurately follow a desired trajectory. Hence, it needs to continuously adjust the actuators to minimize the error in position, ensuring accurate motion control. To achieve this, the PID controller has been implemented using spiking neurons to represent the error signals and components; see Fig. 27(b). However, it relies on supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge and thus is not suitable for the OCL scenario.
@@ IV-B1 Cars Detection
Several works have been developed for low-power autonomous vehicles using SNNs, including studies of cars detection in CarSNN [343], SNN4Agents [324], FastSpiker [344] and [345, 346]. CarSNN and SNN4Agents used an attention window mechanism (i.e., focusing on regions with the highest event density) to process event-based input samples from the N-CARS dataset [347]; see Fig. 28. However, the state-of-the-art [343, 324] require a long training time [344]. Therefore, FastSpiker proposed a methodology to accelerate SNN training while maintaining accuracy through learning rate enhancements [344]. However, these works considered supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, hence they are not suitable for OCL scenario.
@@ IV-B2 Road Lane Detection
Other SNN works for low-power autonomous vehicles include the road line detection. For instance, LaneSNN [348] addressed the imbalance between lane and background classes by employing a novel loss function combining Weighted Binary Cross Entropy (WCE) and Mean Squared Error (MSE); see Fig. 29. Evaluations demonstrated comparable performance to state-of-the-art methods like LaneNet and RefineNet but with significantly lower power consumption. However, it still considered supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, hence it is not suitable for OCL scenario.
@@ IV-B3 Simultaneous Localization and Mapping (SLAM)
Figure 30: DVS-Radar Fusion for SNN-STDP based CL SLAM system; adapted from studies in [349].
SLAM is important for an autonomous vehicle/agent for constructing a map while simultaneously keeping track of its location in an unknown environment. The DVS-Radar Fusion SLAM [349] is the first SNN-based method for enabling CL-based SLAM using drones (Fig. 30), integrating DVS and a Frequency Modulated Continuous Wave (FMCW) radar to encode sensory data on-the-fly into feature descriptors. To provide drones’ velocity information and aid in obstacle detection, a radar-gyroscope odometry method is used for accurate navigation and mapping. SNN output is utilized as feature descriptors for loop closure detection that are fed to a RatSLAM back-end. This process helps in identifying when the drone revisits a previously mapped area. The obstacle modelling is achieved using radar data by detecting objects and integrating this information into its SLAM framework. This work showed the potential of implementing OCL on SNN-based mobile agents.

### v3

@@ 1 FC layers
2 FC layers 68% (6-bit quantized weights) Unsupervised Class-IL 8x weight memory reduction with 4-bit than 32-bit n/a n/a
[356] 4 FC layers 80.0% on FPGA Supervised Task-IL 35.3% flip-flops, 29.3% BRAMs n/a n/a
ImageNet
X-DER[199] EfficientNet-B2 28.19% Supervised Class-IL $2000^{1}$ n/a n/a
SDAF [396] ResNet18 & MLP $33.2\%\pm 0.5$ OCL Class-IL $5000^{1}$ n/a n/a
DVC [395] Resnet18 $19.1\%\pm 0.9$ OCL Class-IL $5000^{1}$ n/a n/a
[223] ResNet18 $39.5\%\pm 0.3$ Supervised Task-IL n/a n/a n/a
SCR [193] ResNet18 & MLP $35.4\%\pm 0.5$ OCL Class-IL $5000^{1}$ n/a n/a
SOR-SNN [361] ResNet-18 ¿55% Supervised Class-IL 0.32M
parameters
n/a n/a
  • 1 Memory buffer size; 2 Approximate time one training epoch took for 24k training patterns; 3 Inference time of an image; Note: Rows with    indicate NCL methods and    indicate DNN-based CL methods.
III-H Trade-Off Analysis
The discussion in this section addresses the key question Q9. Hardware implementation of NCL typically aims at improving performance (speed-up) and computational efficiency (e.g., in terms of power/energy consumption), while offering high accuracy. However, achieving high accuracy with high speed-up and high efficiency in NCL is a challenging task due to the conflicting nature of their requirements. For instance, in order to achieve high accuracy, the network model typically requires more resources (e.g., more neurons and weights) to provide more memory for continually storing new knowledge from learning new tasks without forgetting, thus significantly increasing the model size.
Energy-Accuracy Trade-offs: SpikeDyn [42] demonstrated substantial energy savings with 51% reduction in training energy and 37% in inference energy for 400 neuron network, (see Table XII), while achieving 21% accuracy improvement over baseline method [102] for new tasks and 8% improvement for previously learned tasks. For smaller 200 neuron network, the trade-offs were even more favorable, showing 57% training and 51% inference energy reductions with 23% accuracy improvements. These results indicate that careful architectural optimizations can yield significant energy benefits without accuracy losses.
Memory-Accuracy Trade-offs: Compressed latent replay methods [299, 363] demonstrated memory-accuracy trade-offs, achieving 140× memory reduction (from 22.4 MB to 160 KB) with only 4% accuracy degradation in Sample-IL scenario [299]. In Class-IL scenario, this method achieved $\approx$86accuracy on layer 3 in case of a 1:5 ratio with 320 kB memory usage, compared to naive rehearsal (without compression) requiring 22.4 MB for 89.55% accuracy, representing a 7× memory efficiency improvement with $\approx$3accuracy drop [299]. Replay4NCL [363] revealed that storing compressed latent replays at layer 3 and replaying them with fewer timesteps (i.e., 40 instead of 100) with critical parameters adjustment achieved 90.43% accuracy compared to 86.22% of the baseline [299], with 4.88x speed-up, 20% latent memory saving, and 36.43% energy saving. lpSpikeCon [43] showed that lowering weight precision from 32-bit to 4-bit degrades task-specific accuracy due to information loss, while achieving 8× memory reduction.
Hardware Platform Energy-Latency Trade-offs: Platform-specific performance showed significant latency variations in [42]: 1.71s per inference on Jetson Nano (10W) versus 0.2s on RTX 2080 Ti (250W), indicating that energy-constrained platforms require 8.5× longer processing time but consume 25× less power, providing clear energy-latency trade-off quantification for different deployment scenarios.
Moreover, maintaining high accuracy often requires the model to employ high precision data format, which may not be supported in commodity neuromorphic hardware [330], thereby making it even more difficult to reduce the network size. On the other hand, neuromorphic hardware platforms typically have limited compute and memory resources, which constrains how the network model will be implemented (i.e., mapped and executed) on the hardware fabric. Therefore, a larger model usually requires longer processing latency, and hence higher power/energy consumption. The DSD-SNN model achieved an efficient trade-off between network compactness and CL by dynamically adapting its structure [360]. On CIFAR100, it stabilized at 37.48% of the full network size while maintaining superior accuracy than the non-pruned variant, which rapidly exhausted memory, leading to performance collapse after six tasks. This highlights the effectiveness of pruning in reducing overhead without compromising learning performance. In some cases, a very large model may even need to be split into several parts (e.g., layer-based partition), so that each part can be mapped and executed on the hardware fabric. Consequently, this condition leads to higher processing latency (lower speed-up) and higher power/energy consumption (lower efficiency).
To address this, trade-off analysis is required. This analysis aims at identifying the appropriate network model candidates considering the given hardware constraints (e.g., compute and memory budgets), and then guides the selection on the most suitable one [42]. For instance, if we can accept a slight accuracy degradation for a targeted application, we may be able to decrease the model size significantly through some optimization techniques (such as quantization and pruning), and hence reducing processing latency (higher speed-up) and power/energy consumption (higher efficiency) [43][266].
III-I Summary of NCL
The state-of-the-art methods have performed initial studies for enabling NCL (Table IX). We provide a comparative quantitative analysis (i.e., numerical results) considering design factors (i.e., network complexity) and key evaluation metrics (accuracy, memory footprint, latency, power/energy usage) of the reviewed NCL methods with relevant state-of-the-art DNN-based CL methods (from Section II-G), see Table XI. The table is designed to align according to evaluated datasets (i.e., P-MNIST, Split CIFAR-100, Split-MNIST, and ImageNet) and includes the performance metrics (as reported in the respective studies of each method). Several reviewed NCL methods like SpNCN [303], SOR-SNN [361] and DSD-SNN [360] explicitly compare their method with established DNN-CL baselines, such as EWC [72], SI [68], MAS [153]. For more detailed comparisons, we refer to the respective studies of the NCL methods.
@@ III-I Summary of NCL
We observe that while optimization-based CL methods in DNNs (i.e., GPM [215][223], A-GEM [222] and GSS [115]) report high accuracy on standard benchmarks such as P-MNIST, the NCL method HLOP-SNN[365] achieves superior accuracy (95.15%) with the same CL setting and scenario. Moreover, the NCL method SOR-SNN [361] achieves competitive accuracy (86.65%) on CIFAR-100 with significantly lower memory (0.32M params) than DNN-based CL methods, highlighting the energy-efficient scalability of spiking architectures. On the MNIST, NCL methods such as DSD-SNN [360] and SpikeDyn[42] report comparable or better accuracy with drastic reductions in latency, memory, and energy (e.g., 0.2s, ¡3000KB), demonstrating the suitability of NCL for lightweight and resource-constrained applications. While accuracy remains lower for both domains on ImageNet, NCL methods like SOR-SNN [361] show promising results in parameter efficiency and deployment feasibility. Furthermore, the NCL methods with bio-plausible learning rules (e.g., STDP) facilitate unsupervised learning, which is suitable for OCL scenario. These works mainly employ FC SNN architectures (see Table XII), consisting of input layer and a pair of excitatory-inhibitory layer. As the input data grows in complexity, FC SNNs struggle to capture the important hierarchical features without increasing the network size. Larger architectures often lead to better accuracy but require larger storage sizes, which may exceed resource constraints. This limits the deployable networks on practical hardware platforms, and makes their performance limited to simple datasets (e.g., MNIST) with simple CL scenarios (e.g., Split-MNIST). Meanwhile, NCL methods that employ supervised-based learning rules (e.g., SG-based BP) can achieve acceptable performance on complex datasets (e.g., CIFAR-100 and ImageNet).
However, they are not suitable for OCL scenario, as they require costly labeled data and power-hungry training. Therefore, alternative methods are required especially for enabling energy-efficient SNNs with OCL capabilities, which are beneficial for many real-world application use-cases. Although, we provide the comparison (Table XI), we acknowledge that perfect fairness remains a challenge due to differences in CL scenarios, hyperparameters, underlying model architectures, training paradigms, hardware assumptions, encoding schemes, evaluation protocols, and CL settings (i.e., OCL, supervised or unsupervised). We have emphasized these limitations in our discussion (Section-V). Meanwhile, we ensured that our comparisons are as meaningful and aligned, as possible under current benchmarking constraints. Thus, providing a contextual foundation for assessing the relative strengths and limitations of approaches and the reliability of results.
Replay-based CL in SNNs differs from its ANN counterparts in terms of operation as the ANN-based replay methods like iCaRL [108], DGR [202] rely on storing raw inputs or latent vectors and updating weights using full gradient feedback. In contrast, SNN-based replay must preserve temporally precise spike trains under event-driven constraints, making it dependent on local or SG learning and limited by encoding complexity and latency alignment. While ANN replay offers high gradient precision and plasticity, SNN replay often relies on approximations like compressed or spike-rate encoded replay to conserve power and memory, which may reduce accuracy. Nonetheless, SNN replay methods such as latent replay or STDP-based approaches are significantly more energy-efficient and well-suited for OCL in low-power environments. Additionally, spike noise from stochastic firing and temporal jitter, along with hardware-induced variability such as component mismatch and drift, introduces instability in synaptic updates and membrane dynamics. These factors hinder reproducibility and memory consolidation in SNNs, necessitating robustness-aware learning rules like homeostatic STDP or meta-plasticity, such challenges are absent in conventional hardware.
@@
TABLE XII: Quantitative analysis of the state-of-the-art NCL in terms of network architecture, model size, performance, memory footprint, and power/energy consumption.
@@
Work SNN Architecture Model Size Performance Memory Footprint Power/Energy Consumption
ASP [102] Hierarchical SNN 6400 excitatory Achieved avg. accuracy of 94.2% outperforming n/a n/a
neurons standard STDP
SpikeDyn 2 FC layers (1 input 200 and 400 Improved accuracy by 23% for the new task & $<$3000 KB Avg. reduction of 57%
[42] & 1 excit. layers) excit. neurons 4% for old tasks than ASP for 200-neuron SNN (training) and 51%
@@
2048, 1024 neurons/layer for Split-MNIST and 74% for MNIST-DVS
@@
Replay4NCL 4 Recurrent-FC 200-100-50-20 Achieved Top-1 accuracy of 90.43% on old knowl- 20% reduction in 4.88x latency improvement
[363] layers neurons edge, outperforming the state-of-the-art latent memory usage and 36.43% energy reduction
than [299]
[294] 3 CONV layers 30-250-200 Achieved $92.0\pm 0.1\%$ accuracy for digits in n/a n/a
@@ IV Real-World Application Use-Cases
In this section, we discuss the real-world application use-cases that will benefit from bio-plausible SNNs with OCL capabilities, taxonomy shown in Fig. 31. Additionally, we also highlight the current state-of-the-art for emerging application use-cases, and hence addressing the key question Q10. A summary of real-world application use-cases that will benefit from SNNs with OCL, is presented in Table XIII. Quantitative analysis (i.e., numerical results) of the case studies covered by these works are reported in Table XIV.
Figure 31: A taxonomy of real-world application use-cases covered in this survey, that will benefit from bio-plausible SNNs with OCL capabilities. We have highlighted the main categories (blue blocks), with sub-categories (red blocks), and each of their works shown (green blocks).
@@ IV-A1 Object Recognition for Humanoid Robots
Adaptive Robots need to identify, classify, and locate objects within an image/video. For this, Neural Engineering Framework (NEF) is explored for developing Neural State Machine (NSM) framework leveraging SNNs, managing the allocation of new neurons (Fig. 32), where data may be collected on-the-fly without any labels [397]. A case study on the interactive continual object learning of the iCub robot was presented in [397]; see Table XIII. Experiments in the Gazebo robotic simulator [398] mimicking event-based camera conditions, and implementations on Intel’s Loihi neuromorphic chip [328], demonstrated $96.55\pm 2.02\%$ accuracy in learning object representations within three epochs, adapting efficiently to new objects [397]; see Table XIV. CL assessments showed rare confusions ($<20\%$) after the first epoch. The model allocated neurons dynamically based on object complexity, optimizing resource use. The continual classifier was benchmarked against conventional online classifiers on a CPU, showing up to 300× better energy efficiency for learning and 150× for inference [397]. These findings highlight SNNs’ potential for real-world deployment in embedded AI systems, with future extensions to physical robotic platforms. However, it relies on supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, thus it is not suitable for OCL scenario.
@@
IV-A2 Robotic Control Systems
The study [399] explored NEF for developing neuromorphic algorithms for Inverse Kinematics (IK) and Proportional-Integral-Derivative (PID) control systems, supporting online learning of control signals, using Prescribed Error Sensitivity (PES). Fig. 33(a) shows the model schematic for neuromorphic IK with online learning, whose target is to efficiently compute the joint angles required to place the robots’ end-effector at the desired position. Meanwhile, the PID controller aims to ensure that the robots’ actuators accurately follow a desired trajectory. Hence, it needs to continuously adjust the actuators to minimize the error in position, ensuring accurate motion control. To achieve this, the PID controller has been implemented using spiking neurons to represent the error signals and components; see Fig. 33(b). This work presented a case study on the control of a 6-degrees of freedom robotic arm; see Table XIII. The implementation was based on Nengo [400, 401] a Python based neural compiler, that translates high-level descriptions to low-level neural models and Intel Loihi neuromorphic chip [328], offering high performing and energy-efficient neuromorphic control; see Table XIV. However, it relies on supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge and thus is not suitable for the OCL scenario.
@@ IV-A2 Robotic Control Systems
Similarly, a recent work in [402] used NEF within the Nengo simulator and MATLAB R2022b to implement a SNN-based PID controller for robotic arm trajectory tracking. The controller was evaluated on a case study of simulated 3-DoF robotic arm; see Table XIII. It reported achieving a 6% improvement in the Integral of Time-weighted Absolute Error (ITAE) and a 30% reduction in Root Mean Square Error (RMSE) compared to the conventional PID and fuzzy controllers; see Table XIV.
@@ IV-B1 Cars Detection
Several works have been developed for low-power autonomous vehicles using SNNs, including studies of cars detection in CarSNN [403], SNN4Agents [381], FastSpiker [404] and [405, 406]. CarSNN and SNN4Agents used an attention window mechanism (i.e., focusing on regions with the highest event density) to process event-based input samples from the N-CARS dataset [341]; see Fig. 34. For the cars vs. background classification the SNN model was implemented on the Intel Loihi neuromorphic chip [328] (see Table XIII). The neuromorphic hardware implementation had maximum 0.72 ms of latency for every sample, and consumed only 310 mW power with an accuracy of 83% [403] (see Table XIV). However, the state-of-the-art [403, 381] require a long training time [404]. Therefore, FastSpiker proposed a methodology to accelerate SNN training while maintaining accuracy through learning rate enhancements [404]. However, these works considered supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, hence they are not suitable for OCL scenario.
@@ IV-B2 Road Lane Detection
Other SNN works for low-power autonomous vehicles include the road line detection. For instance, LaneSNN [407] addressed the imbalance between lane and background classes by employing a novel loss function combining Weighted Binary Cross Entropy (WCE) and MSE; see Fig. 35. For detecting the lanes marked on the streets using the event-based camera input, the SNNs training and implementation was based on PyTorch library [319] and the Intel Loihi neuromorphic chip [328]; see Table XIII. Evaluations demonstrated maximum latency of less than 8 ms, power consumption of about 1 W during the classification of a single image, and online IoU equal to 0.623, thereby making it superior to the state-of-the-art techniques like LaneNet and RefineNet in terms of performance and power efficiency; see Table XIV. However, it still considered supervised settings which require costly labeled data, power-hungry training, and long training time to update its knowledge, hence it is not suitable for OCL scenario.
@@ IV-B3 Simultaneous Localization and Mapping (SLAM)
Figure 36: a) Setup on a drone in a UWB-equipped warehouse for ground truth localization, b) Event and radar data are processed by an STDP-trained SNN for loop closure detection in SLAM, with radar also modeling obstacles (black dots), and c) DVS-Radar Fusion for SNN-STDP based CL SLAM system [408].
SLAM is important for an autonomous vehicle/agent for constructing a map while simultaneously keeping track of its location in an unknown environment. The DVS-Radar Fusion SLAM [408] is the first SNN-based method for enabling CL-based SLAM using drones (Fig. 36), integrating DVS and a Frequency Modulated Continuous Wave (FMCW) radar to encode sensory data on-the-fly into feature descriptors. To provide drones’ velocity information and aid in obstacle detection, a radar-gyroscope odometry method is used for accurate navigation and mapping. SNN output is utilized as feature descriptors for loop closure detection that are fed to a RatSLAM back-end. This process helps in identifying when the drone revisits a previously mapped area. The obstacle modelling is achieved using radar data by detecting objects and integrating this information into its SLAM framework. DVS-Radar fusion setup outperformed RGB-based methods achieving lower $MAE_{L}$ and $MAE_{M}$ values of 0.51 & 0.17 for drone flight sequence 1, and 0.81 & 0.45 for sequence 2; see Table XIII and XIV. This work showed the potential of implementing OCL on SNN-based mobile agents.
TABLE XIII: Summary of recent progress in real-world NCL application use-cases that will benefit from SNNs with OCL capabilities, with their implementation details and case studies.
Application Use- Work Case CL Dataset CL Learning Neuron HW
Domains cases (Year) Study Setting Approach Rule Model
Object [397] Interactive Supervised Custom Representation- Modified LIF Loihi
Recognition (2022) continual object DVS dataset based Hebbian
Adaptive learning in iCub robot
Robots Control of Custom
Robotic Arm [399] (2021) 6-DoF, Supervised dataset from Representation- PES LIF Loihi
Control [402] (2024) 3-DoF trajectory robotic arms’ based
tracking operations
Cars [403] (2021) Cars vs. STBP LIF Loihi
Detection [405] (2022) background Supervised N-CARS - SG LIF GPU
[381] (2024) classification STBP LIF GPU
Road Lane [407] Detection of
Autonomous Detection (2022) lanes marked Supervised DET [352] - STBP LIF Loihi
Vehicles/ on the streets
Agents [408] CL SLAM Real-time Representation-
SLAM (2023) fusing DVS & Unsupervised data from based STDP LIF n/a
radar on drone sensors
People [409] CL of people KUL- ReckOn
Detection (2024) detection from Semi-supervised UAVSAFE Representation- STDP LIF $chip^{*}$
event-camera [353] based [410]
mounted on drone
Robotic [393] Tracking & CIFAR-10,
Navigation (2024) obstacle avoidance Hybrid self-built - - LIF PAIBoard
in robot dog [411] dataset
  • * Potential future work.
TABLE XIV: Quantitative analysis of case studies covered by the reviewed works in Section IV.
Work Performance OCL Latency Memory Power/Energy
Capability Footprint Savings
[397] 96.55±2.02% new object accuracy within $\times$ n/a n/a 300× better for training
three epochs, ¡20% confusions after first epoch & 150× for inference
[399] Loihi accuracy outperformed the simulation $\times$ Loihi implementation converged faster n/a Improved
across different learning rates to the target than the simulated model
 [402] 6% improvement in ITAE and 30% reduction $\times$ n/a n/a n/a
in RMSE compared to traditional PID controller
 [403] 83% accuracy $\times$ 0.72 ms n/a 310 mW
 [407] online IoU 0.623 $\times$ 8 ms n/a $1W^{1}$
 [408] $MAE_{L}$ of 0.51 & $MAE_{M}$ of 0.17 for drone flight
sequence 1, $MAE_{L}$ of 0.81 & $MAE_{M}$ of 0.45 for n/a n/a n/a
sequence 2
[409] Peak $F_{1}$ score 19% higher compared to same- n/a 514 kB n/a
size CNN
[393] 90.2% accuracy n/a n/a $energy^{2}$ $791FPS/W$,
$power^{3}$ $12.8W$
  • 1 For single image classification; 2 For image classification tasks; 3 For tracking and obstacle avoidance task.